OpenStudy (anonymous):

In coming to a stop, a car leaves skid marks 96.0 m long on the highway. Assuming a deceleration of 6.40 m/s2, estimate the speed of the car just before braking.

OpenStudy (anonymous):

2as=v^2-u^2 2*6.4*96=v^2-0 Hope that helps. :)